(no subject)

arenton@uvahef.phys.virginia.edu
Sat, 29 May 1999 01:03:58 -0400

Shift Summary for evening shift, 28/may/99

On Shift: M.Arenton and S.Taegar plus about a score of
others.

Hours of beam delivered at all intensities: 2.5
Hours of beam delivered at nominal intensity:
Hours of data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 2.5

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 13195 first run with muons!!
Run 13197 after -3 ns trigger timing change, first
run for DC plateau.

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:
to turn CsI low volts back on after air conditioning repair

========================================================================
Description of Shift:

NM3,4 interlocked just before start of shift.

CsI low voltage off due to air conditioning problems,
finally about 18:00 repair person arrived and fixed the problem.
Made an access to turn CsI low voltage back on and install
cables for new DC-or trigger sources.

20:10 first run with CsI back on shows that channel 206 OK
, so DPMT swap earlier today fixed problem.

21:28, first weak pulse of muons!

22:17 start run 13195 with beam_muon, prescale set to 1.

22:56 change trigger timing by -3 ns. start run 13197.
early looks at CsI indicate a similar timing shift.

23:00 beginning series of runs for DC plateau.

23;45 lower sweeper currents to NM2S2 500A, NM2S3 100A, this
increases rate, earlier rates were reasonably the same as
in 1997, SEM has been below 1E11. Note that channel 13 display
has little useful information, it says no-beam for all
accelerators whether of not there really is beam.